3.1The following data (R billions) are for a closed economy (Ga-Rankuwa Republic)
over 2019.
• Total Output: R 30
• Total Income: R 30
• Consumption: R 22
• Government Spending: R 6
• Total tax revenue: R 5.5
• Transfer payments: R 1.5
Use the above data to calculate:
a. disposable income, savings and net taxes.
b. Is the government running a budget surplus, budget deficit, or balanced
budget? Demonstrate.
c. Determine planned investment. (note that investment is equal to the sum of
private and public saving)
1
Expert's answer
2020-07-01T15:37:08-0400

1a)Disposable income =Total income - Tax

= R. 30-R5.5=R 24.5

b) Savings =Total Output - Consumption - Government Spending

=30-22-6=R2

C) Net Taxes =Government Spending

=R. 6

2)The government is running a balanced budget since the government revenue is equal to the the government spending . i.e Total Government Spending(government spending +transfer payments) =R7.5 and Government Revenue(Taxes +savings) =R 7.5

3)planned investment =Public savings + private savings.

(Savings +transfer payments)

=(2+1.5)=R 3.5